OpenStudy (anonymous):

Angelina washes 13 1/2 windows in 2/3 hour. At this rate, how many windows can she wash in 1 hour?

OpenStudy (jdoe0001):

\(\begin{array}{ccllll} windows&hours \\\hline\\ 13\frac{1}{2}&\cfrac{2}{3}\\ x&1 \end{array}\implies \cfrac{13\frac{1}{2}}{x}=\cfrac{\frac{2}{3}}{1}\) solve for "x"

OpenStudy (anonymous):

Ok!

OpenStudy (jdoe0001):

so.. convert the 13 1/2 to "improper" fraction, that is, just not a mixed one and find "x" :)

OpenStudy (jdoe0001):

recall that \(\Large \cfrac{\frac{a}{b}}{\frac{c}{{\color{blue}{ d}}}}\implies \cfrac{a}{b}\cdot \cfrac{{\color{blue}{ d}}}{c}\)

OpenStudy (anonymous):

I got 9 windows!

OpenStudy (jdoe0001):

hmm what did you get for 13 1/2 as an improper fraction anyway?

OpenStudy (anonymous):

i got 27/2

OpenStudy (jdoe0001):

well.. .yes... can't be 9 though notice she does 13 and half windows is just 2/3 of one hour how can in 1 hour does 9? kinda like the longer the she works, the less she does =) one sec, lemme post something

OpenStudy (anonymous):

ok

OpenStudy (jdoe0001):

\(\begin{array}{ccllll} windows&hours \\\hline\\ 13\frac{1}{2}&\cfrac{2}{3}\\ x&1 \end{array}\implies \cfrac{13\frac{1}{2}}{x}=\cfrac{\frac{2}{3}}{1}\implies \cfrac{\frac{27}{2}}{\frac{2}{3}}=x \\ \quad \\ \cfrac{27}{2}\cdot \cfrac{3}{2}=x\implies \cfrac{27\cdot 3}{2\cdot 2}=x\implies \cfrac{81}{4}=x \\ \quad \\ 20\frac{1}{4}=x\)

OpenStudy (anonymous):

but doesnt it have to be less?

OpenStudy (anonymous):

oh, wait i understand now!!!!